OpenStudy (anonymous):

A man earns $9.24 for each of the first 36 hours he works in one week and $13.86 in overtime pay for each additional hour he works in the same week. How many hours of overtime did he work if he made $512.82 that week? What was his overtime hours? Enter your answer as a whole or decimal number.

OpenStudy (anonymous):

9,24+13,86(x-36)=512,82

OpenStudy (anonymous):

72,33 hours :)

OpenStudy (anonymous):

so the answer is 72,33 overtime hours

OpenStudy (johnweldon1993):

well if he works 36 hours...at 9.24 each of those hours...thats 36 times 9.24...thats 332.64 the total amount he made is 512.82 512.82-332.64 = 180.18 to find how many overtime hours he worked.... 180.18 / 13.86 this is 13 Am i wrong?

OpenStudy (anonymous):

well some one else said 72,33 hours

OpenStudy (johnweldon1993):

well if he works 72.33 overtime hours....thats 72.33 times 13.86 in overtime pay so that equals 1002.5 dollars...he didnt make anywhere near that...so i would go with what my explanation said.....but please anyone correct me if I'm wrong

OpenStudy (anonymous):

so you said his overtime is 13

OpenStudy (johnweldon1993):

right because if he works 13 hours of overtime...our equation would look like 9.24n + 13.86o = 512.82 ****n means hours of normal pay.....****o means overtime hours so 9.24(36) + 13.86o = 512.82 332.64 + 13.86o = 512.82 13.68o =180.18 o = 13 so 13 overtime hours

OpenStudy (anonymous):

thank you 13 was right

OpenStudy (johnweldon1993):

no problem :)

OpenStudy (anonymous):

can you help me with my next question

OpenStudy (johnweldon1993):

if i can sure :)